# Transversal Hamilton cycles in digraph collections

### v1

@@ 3 Absorbing
In this case, we prove $|Z_{j}^{\rm II}(i,vu)|\geq 2^{-7}\delta^{2}n^{4}$. By Lemma 3.3 (iii), we have either $d_{D_{i}}^{+}(v,W_{i}^{1})\geq(\frac{1}{2}-3\epsilon)n$ or $d_{D_{i}}^{+}(v,W_{i}^{3})\geq(\frac{1}{2}-3\epsilon)n$. Without loss of generality, we assume that the former case holds since the latter case can be proved similarly. Note that $D_{i}$ and $D_{j}$ are $\delta$-crossing. Then $|W_{i}^{1}\cap W_{j}^{1}|\geq\frac{\delta n}{4}$ and $|W_{i}^{1}\cap W_{j}^{3}|\geq\frac{\delta n}{4}$. Therefore,
$d_{i}^{+}(v,W_{j}^{1})\geq|W_{i}^{1}\cap W_{j}^{1}|-\left(|W_{i}^{1}|-d_{i}^{+}(v,W_{i}^{1})\right)\geq\frac{\delta n}{4}-2\epsilon n\geq\frac{\delta n}{5}\ (\text{since}\ \epsilon\ll\delta).$
@@
Notice that $|L_{i}|=2\epsilon n$ when $D_{i}$ is $\epsilon$-extremal. Hence $|T_{i}|\frac{\delta n}{2}\leq 2\epsilon n^{2}$, that is, $|T_{i}|\leq\frac{\sqrt{\epsilon}}{2}n$ since $\epsilon\ll\delta$.
Let $\overline{T_{i}}:=V\setminus T_{i}$. Then $|\overline{T_{i}}|\geq(1-\frac{\sqrt{\epsilon}}{2})n$. For each $i\in V(H)$ and $u\in\overline{T_{i}}$, we have $|N_{H}(i)\cap\mathcal{C}_{u}|\geq\frac{\delta n}{2}$. Now we independently and randomly select vertices from $V(H)$ with probability $\kappa:=\frac{\lambda}{14}$ to obtain a color set $\mathcal{U}$. By using Chernoff’s bound, with high probability the following hold:
@@
Given $i\in\overline{\mathcal{U}}$ and $u\in\overline{T_{i}}$, choose $j\in\mathcal{C}_{u}\cap N_{H}(i,\mathcal{U})$. Then $\frac{j+1}{3}\in[t]$ and the number of such choices of $j$ is at least $\frac{\delta\kappa n}{4}\geq\frac{\delta t}{8}$ by (**∗ ∗). Let $v$ be a $D_{i}$-good vertex. It follows from Claim 3.10 that $|Z_{j}^{\rm K}(i,vu)|\geq 2^{-7}\delta n^{4}$ for some ${\rm K}\in\{{\rm I,II}\}$.
@@
[…]
@@
Relabel indices such that $I=:[s]$ with $s\geq(1-2^{-16}\delta^{2})t$. We are to connect those disjoint rainbow directed paths $P^{1},\ldots,P^{s}$ into a directed absorbing cycle $C$. Notice that $\overline{\mathcal{U}}\cap{\rm col}(\bigcup_{i\in I}P^{i})=\emptyset$. We first connect $P^{1}$ and $P^{2}$ into a single rainbow directed path $P^{1}v_{4}^{1}xzyv_{1}^{2}P^{2}$ or $P^{1}v_{4}^{1}xzv_{1}^{2}P^{2}$. For this, choose distinct colors $c_{1},c_{2},c_{3},c_{4}\in\overline{\mathcal{U}}$ such that $c_{1}c_{2}\in E(H)$. Clearly, there are at least $|\overline{\mathcal{U}}|\geq\frac{\delta n}{2}$ choices for $c_{1}$, and given this at least $\delta(H[\overline{\mathcal{U}}])\geq\frac{\delta n}{2}$ choices for $c_{2}$, and at least $|\overline{\mathcal{U}}|-3>\frac{\delta n}{3}$ choices for each of $c_{3},c_{4}$. Next, choose an unused $D_{c_{1}}$-good vertex $x\in N^{+}_{{c_{3}}}(v_{4}^{1})$ and an unused $D_{c_{2}}$-good vertex $y\in N^{-}_{{c_{4}}}(v_{1}^{2})$.
@@
  • there exists a rainbow directed path $P^{1}v_{4}^{1}xyzv_{1}^{2}P^{2}$ with colors ${\rm col}(P^{1}\cup P^{2})\cup\{c_{1},c_{2},c_{3},c_{4}\}$ and $z\in N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)$, each of $x,y,z$ has at least $\frac{\delta n}{10}$ choices;
@@
[…]
@@
[…]
If $|W_{c_{1}}^{2}\cap N^{+}_{{c_{3}}}(v_{4}^{1})|\geq\delta n$ and $|W_{c_{2}}^{3}\cap N^{-}_{{c_{4}}}(v_{1}^{2})|\geq\delta n$, then choose $x\in W_{c_{1}}^{2}\cap N^{+}_{{c_{3}}}(v_{4}^{1})$ and $y\in W_{c_{2}}^{3}\cap N^{-}_{{c_{4}}}(v_{1}^{2})$. Hence, there are at least $\delta n-7e(M)-2\epsilon n\geq\frac{\delta n}{2}$ choices for each of $x$ and $y$. Furthermore, $|N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)|\geq|W_{c_{1}}^{3}\cap W_{c_{2}}^{2}|-4\epsilon n\geq\frac{\delta n}{5}$, as desired.
If $|W_{c_{1}}^{2}\cap N^{+}_{{c_{3}}}(v_{4}^{1})|<\delta n$, then $|W_{c_{1}}^{4}\cap N^{+}_{{c_{3}}}(v_{4}^{1})|\geq(\frac{1}{2}-\mu-\delta-2\epsilon)n$. Choose $x\in W_{c_{1}}^{4}\cap N^{+}_{{c_{3}}}(v_{4}^{1})$. Then $d_{{c_{1}}}^{+}(x,W_{c_{1}}^{1})\geq|W_{c_{1}}^{1}|-2\epsilon n$. Notice that either $|W_{c_{2}}^{1}\cap N^{-}_{{c_{4}}}(v_{1}^{2})|\geq\delta n$ or $|W_{c_{2}}^{3}\cap N^{-}_{{c_{4}}}(v_{1}^{2})|\geq\delta n$. We may assume the former case holds. Choose $y\in W_{c_{2}}^{1}\cap N^{-}_{{c_{4}}}(v_{1}^{2})$. Then $d_{{c_{2}}}^{-}(y,W_{c_{2}}^{4})\geq|W_{c_{2}}^{4}|-2\epsilon n$. If $|W_{c_{1}}^{1}\cap W_{c_{2}}^{4}|\geq\delta n$, then $|N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)|\geq|W_{c_{1}}^{1}\cap W_{c_{2}}^{4}|-4\epsilon n\geq\frac{\delta n}{5}$, as desired. If $|W_{c_{1}}^{1}\cap W_{c_{2}}^{4}|<\delta n$ and $|W_{c_{2}}^{3}\cap N^{-}_{{c_{4}}}(v_{1}^{2})|\geq\frac{\delta n}{9}$, then $|W_{c_{1}}^{1}\cap W_{c_{2}}^{2}|\geq(\frac{1}{2}-\delta-3\epsilon)n$ and choose $y\in W_{c_{2}}^{3}\cap N^{-}_{{c_{4}}}(v_{1}^{2})$. Similarly, $|N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)|\geq|W_{c_{1}}^{1}\cap W_{c_{2}}^{2}|-4\epsilon n\geq\frac{\delta n}{5}$, as desired. If $|W_{c_{1}}^{1}\cap W_{c_{2}}^{4}|<\delta n$ and $|W_{c_{2}}^{3}\cap N^{-}_{{c_{4}}}(v_{1}^{2})|<\frac{\delta n}{9}$, then $d^{-}_{{c_{4}}}(v_{1}^{2},W_{c_{2}}^{1})\geq|W_{c_{2}}^{1}|-(2\epsilon+\frac{\delta}{9})n$. Therefore,
$|N_{{c_{1}}}^{+}(x)\cap N_{{c_{4}}}^{-}(v_{1}^{2})|\geq|N_{{c_{1}}}^{+}(x,W_{c_{1}}^{1})\cap N_{{c_{4}}}^{-}(v_{1}^{2},W_{c_{2}}^{1})|\geq|W_{c_{1}}^{1}\cap W_{c_{2}}^{1}|-\left(4\epsilon+\frac{\delta}{9}\right)n\geq\frac{\delta n}{9}.$
Now, we connect $P^{1}$ and $P^{2}$ by using a directed rainbow path $P^{1}v_{4}^{1}xyv_{1}^{2}P^{2}$ with colors ${\rm col}(P^{1}\cup P^{2})\cup\{c_{1},c_{3},c_{4}\}$, as desired.
@@
[…]
Now, we consider $|Z_{c_{1}}\cap W_{c_{2}}|<\frac{\delta n}{4}$. Without loss of generality, assume that $Z=A$ and $W=W^{2}$. It follows that $|A_{c_{1}}\cap W_{c_{2}}^{4}|\geq|W_{c_{2}}^{4}|-(\frac{\delta}{4}+2\epsilon)n$. This implies $|A_{c_{1}}\cap W_{c_{2}}^{3}|\geq|C_{c_{2}}^{4}|-(\frac{\delta}{4}+2\epsilon)n\geq\frac{\delta n}{2}$ since $|C^{2}_{4}|\geq\delta n$. If $|W_{c_{2}}^{1}\cap N_{{c_{4}}}^{-}(v_{2}^{1})|\geq\frac{\delta n}{9}$, then choose $y\in W_{c_{2}}^{1}\cap N_{{c_{4}}}^{-}(v_{2}^{1})$. Hence $|N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)|\geq|A_{c_{1}}\cap W_{c_{2}}^{4}|-3\epsilon n\geq\frac{\delta n}{5}$. If $|W_{c_{2}}^{1}\cap N_{{c_{4}}}^{-}(v_{2}^{1})|<\frac{\delta n}{9}$, then $|W_{c_{2}}^{3}\cap N_{{c_{4}}}^{-}(v_{2}^{1})|\geq|W_{c_{2}}^{3}|-(\frac{\delta}{9}+2\epsilon)n$. Thus, $|N_{{c_{1}}}^{+}(x)\cap N_{{c_{4}}}^{-}(v_{2}^{1})|\geq|A_{c_{1}}\cap W_{c_{2}}^{3}|-(3\epsilon+\frac{\delta}{9})n\geq\frac{\delta n}{9}$. Now, we connect $P^{1}$ and $P^{2}$ by using a directed rainbow path $P^{1}v_{4}^{1}xyv_{1}^{2}P^{2}$ with colors ${\rm col}(P^{1}\cup P^{2})\cup\{c_{1},c_{3},c_{4}\}$, as desired.
@@
In view of Claim 3.11, there are at least $\frac{\delta n}{10}$ choices for each of $c_{1},c_{2},c_{3},c_{4},x,y,(z)$ given any previous choices. Thus, we obtain the desired directed rainbow path $P^{1}v_{4}^{1}xzyv_{1}^{2}P^{2}$ or $P^{1}v_{4}^{1}xyv_{1}^{2}P^{2}$. Applying the above process for each pair $(P^{j},P^{j+1})$ with $j\in[s]$ where $P^{s+1}:=P^{1}$. Every time we use at most four unused colors in $\overline{\mathcal{U}}$ and at most three unused vertices. Therefore, so for each choice, at most $4s\leq 4t\leq 8\kappa n\ll\frac{\delta n}{10}$ colors are forbidden and $3s\ll\frac{\delta n}{10}$ vertices are forbidden. Thus we obtain a directed rainbow cycle $C$ of length at most $7s$ which contains every $P^{j}$ with $j\in[s]$ as a segment.

### v2

@@ 3 Absorbing
In this case, we prove $|Z_{j}^{\rm II}(i,vu)|\geq 2^{-7}{{\delta}}n^{4}$. By Lemma 3.2 (iii), we have either $d_{D_{i}}^{+}(v,W_{i}^{1})\geq(\frac{1}{2}-3\epsilon)n$ or $d_{D_{i}}^{+}(v,W_{i}^{3})\geq(\frac{1}{2}-3\epsilon)n$. Without loss of generality, we assume that the former case holds since the latter case can be proved similarly. Note that $D_{i}$ and $D_{j}$ are $\delta$-crossing. Then $|W_{i}^{1}\cap W_{j}^{1}|\geq\frac{\delta n}{4}$ and $|W_{i}^{1}\cap W_{j}^{3}|\geq\frac{\delta n}{4}$. Therefore,
$d_{i}^{+}(v,W_{j}^{1})\geq|W_{i}^{1}\cap W_{j}^{1}|-\left(|W_{i}^{1}|-d_{i}^{+}(v,W_{i}^{1})\right)\geq\frac{\delta n}{4}-2\epsilon n\geq\frac{\delta n}{5}\ (\text{since}\ \epsilon\leq\delta^{2}).$
@@
Notice that $|L_{i}|=2\epsilon n$ when $D_{i}$ is $\epsilon$-extremal. Hence $|T_{i}|\frac{\delta n}{2}\leq 2\epsilon n^{2}$, that is, $|T_{i}|\leq\frac{\sqrt{\epsilon}}{2}n$ since $8\epsilon^{1/2}\leq\delta$.
For each $i\in V(H)$, let $\overline{T_{i}}:=V\setminus T_{i}$. Then $|\overline{T_{i}}|\geq(1-\frac{\sqrt{\epsilon}}{2})n$. For each $u\in\overline{T_{i}}$ with $i\in V(H)$, one has $|N_{H}(i)\cap\mathcal{C}_{u}|\geq\frac{\delta n}{2}$. Now we independently and randomly select vertices from $V(H)$ with probability $\kappa:=\frac{\lambda}{14}$ to obtain a color set $\mathcal{U}$. Using Chernoff’s bound, with high probability the following hold:
@@
Given $i\in\overline{\mathcal{U}}$ and $u\in\overline{T_{i}}$, choose $j\in\mathcal{C}_{u}\cap N_{H}(i,\mathcal{U})$. Then $\frac{j+1}{3}\in[t]$ and the number of such choices of $j$ is at least $\frac{\delta\kappa n}{4}\geq\frac{\delta t}{8}$ by (**). Let $v$ be a $D_{i}$-good vertex. It follows from Claim 3.10 that $|Z_{j}^{\rm K}(i,vu)|\geq 2^{-7}\delta n^{4}$ for some ${\rm K}\in\{{\rm I,II}\}$.
@@
$\mathbf{Z}:=\left\{Z(i,vu):=\bigcup_{j\in\mathcal{C}_{u}\cap N_{H}(i,\mathcal{U})}Z_{j}^{\rm K}(i,vu):i\in\overline{\mathcal{U}},\ u\in\overline{T_{i}}\ \text{and}\ v\ \text{is}\ D_{i}\text{-good}\right\},$
@@
Relabel indices such that $I=:[s]$ with $s\geq(1-2^{-16}\delta^{2})t$. We are to connect those disjoint rainbow directed paths $P^{1},\ldots,P^{s}$ into a directed absorbing cycle $C$. Notice that $\overline{\mathcal{U}}\cap{\rm col}(\bigcup_{i\in I}P^{i})=\emptyset$. We first connect $P^{1}$ and $P^{2}$ into a single rainbow directed path $P^{1}v_{4}^{1}xzyv_{1}^{2}P^{2}$ or $P^{1}v_{4}^{1}xzv_{1}^{2}P^{2}$. For this, choose distinct colors $c_{1},c_{2},c_{3},c_{4}\in\overline{\mathcal{U}}$ such that $c_{1}c_{2}\in E(H)$. Clearly, there are at least $|\overline{\mathcal{U}}|\geq\frac{\delta n}{2}$ choices for $c_{1}$, and given this at least $\delta(H[\overline{\mathcal{U}}])\geq\frac{\delta n}{2}$ choices for $c_{2}$, and at least $|\overline{\mathcal{U}}|-3>\frac{\delta n}{3}$ choices for each of $c_{3},c_{4}$. Next, choose an unused $D_{c_{1}}$-good vertex $x\in N^{+}_{{c_{3}}}(v_{4}^{1})$ and an unused $D_{c_{2}}$-good vertex $y\in N^{-}_{{c_{4}}}(v_{1}^{2})$. The following claim considers the number of choices for $x,y$.
@@
  • there exists a rainbow directed path $P^{1}v_{4}^{1}xzyv_{1}^{2}P^{2}$ with colors ${\rm col}(P^{1}\cup P^{2})\cup\{c_{1},c_{2},c_{3},c_{4}\}$ and $z\in N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)$, each of $x,y,z$ has at least $\frac{\delta n}{10}$ choices;
@@
In this case, $D_{c_{1}}$ and $D_{c_{2}}$ are either $(\epsilon,{\rm EC}1)$-extremal or $(\epsilon,{\rm EC}2)$-extremal. Then $x\in A_{c_{1}}\cup B_{c_{1}}$ and $y\in A_{c_{2}}\cup B_{c_{2}}$. Clearly, each of $x,y$ has at least $\frac{n}{4}$ choices. Recall that $D_{c_{1}}$ and $D_{c_{2}}$ are $\delta$-crossing. Then, $|X_{c_{1}}\cap Y_{c_{2}}|\geq\frac{\delta n}{4}$ whenever $X,Y\in\{A,B\}$. By Lemma 3.2, there are $Z,W\in\{A,B\}$ such that $d_{{c_{1}}}^{+}(x,Z_{c_{1}})\geq|Z_{c_{1}}|-\epsilon n$ and $d_{{c_{2}}}^{-}(y,W_{c_{2}})\geq|W_{c_{2}}|-\epsilon n$. Choose $z\in N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)$ and the number of choices for $z$ is at least
@@
In this case, both $D_{c_{1}}$ and $D_{c_{2}}$ are $(\epsilon,{\rm EC}3)$-extremal. Then $|W_{c_{1}}^{i}\cap W_{c_{2}}^{j}|\geq\frac{\delta n}{4}$ for $i,j\in\{1,3\}$. Notice that either $|W_{c_{1}}^{1}\cap W_{c_{2}}^{2}|\geq\frac{n}{5}$ or $|W_{c_{1}}^{3}\cap W_{c_{2}}^{2}|\geq\frac{n}{5}$ holds. Without loss of generality, assume the latter case hold. Choose $z\in N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)$. Next, we consider the size of $|N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)|$.
If $|W_{c_{1}}^{2}\cap N^{+}_{{c_{3}}}(v_{4}^{1})|\geq\delta n$ and $|W_{c_{2}}^{3}\cap N^{-}_{{c_{4}}}(v_{1}^{2})|\geq\delta n$, then choose $x\in W_{c_{1}}^{2}\cap N^{+}_{{c_{3}}}(v_{4}^{1})$ and $y\in W_{c_{2}}^{3}\cap N^{-}_{{c_{4}}}(v_{1}^{2})$. Hence, there are at least $\frac{\delta n}{2}$ choices for each of $x$ and $y$. Furthermore, $|N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)|\geq|W_{c_{1}}^{3}\cap W_{c_{2}}^{2}|-4\epsilon n\geq\frac{\delta n}{5}$, as desired.
If $|W_{c_{1}}^{2}\cap N^{+}_{{c_{3}}}(v_{4}^{1})|<\delta n$, then $|W_{c_{1}}^{4}\cap N^{+}_{{c_{3}}}(v_{4}^{1})|\geq(\frac{1}{2}-\mu-\delta-2\epsilon)n$. Choose $x\in W_{c_{1}}^{4}\cap N^{+}_{{c_{3}}}(v_{4}^{1})$. Then $d_{{c_{1}}}^{+}(x,W_{c_{1}}^{1})\geq|W_{c_{1}}^{1}|-2\epsilon n$. Notice that either $|W_{c_{2}}^{1}\cap N^{-}_{{c_{4}}}(v_{1}^{2})|\geq\delta n$ or $|W_{c_{2}}^{3}\cap N^{-}_{{c_{4}}}(v_{1}^{2})|\geq\delta n$. We may assume the former case holds. Choose $y\in W_{c_{2}}^{1}\cap N^{-}_{{c_{4}}}(v_{1}^{2})$. Then $d_{{c_{2}}}^{-}(y,W_{c_{2}}^{4})\geq|W_{c_{2}}^{4}|-2\epsilon n$. If $|W_{c_{1}}^{1}\cap W_{c_{2}}^{4}|\geq\delta n$, then $|N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)|\geq|W_{c_{1}}^{1}\cap W_{c_{2}}^{4}|-4\epsilon n\geq\frac{\delta n}{5}$, as desired. If $|W_{c_{1}}^{1}\cap W_{c_{2}}^{4}|<\delta n$ and $|W_{c_{2}}^{3}\cap N^{-}_{{c_{4}}}(v_{1}^{2})|\geq\frac{\delta n}{9}$, then $|W_{c_{1}}^{1}\cap W_{c_{2}}^{2}|\geq(\frac{1}{2}-\delta-3\epsilon)n$ and choose $y\in W_{c_{2}}^{3}\cap N^{-}_{{c_{4}}}(v_{1}^{2})$. Similarly, $|N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)|\geq|W_{c_{1}}^{1}\cap W_{c_{2}}^{2}|-4\epsilon n\geq\frac{\delta n}{5}$, as desired. If $|W_{c_{1}}^{1}\cap W_{c_{2}}^{4}|<\delta n$ and $|W_{c_{2}}^{3}\cap N^{-}_{{c_{4}}}(v_{1}^{2})|<\frac{\delta n}{9}$, then $d^{-}_{{c_{4}}}(v_{1}^{2},W_{c_{2}}^{1})\geq|W_{c_{2}}^{1}|-(\mu+2\epsilon+\frac{\delta}{9})n$. Therefore,
$|N_{{c_{1}}}^{+}(x)\cap N_{{c_{4}}}^{-}(v_{1}^{2})|\geq|N_{{c_{1}}}^{+}(x,W_{c_{1}}^{1})\cap N_{{c_{4}}}^{-}(v_{1}^{2},W_{c_{2}}^{1})|\geq|W_{c_{1}}^{1}\cap W_{c_{2}}^{1}|-\left(\mu+4\epsilon+\frac{\delta}{9}\right)n\geq\frac{\delta n}{9}.$
Now, we connect $P^{1}$ and $P^{2}$ by using a rainbow directed path $P^{1}v_{4}^{1}xyv_{1}^{2}P^{2}$ with colors ${\rm col}(P^{1}\cup P^{2})\cup\{c_{1},c_{3},c_{4}\}$, as desired.
@@
In this case, $D_{c_{1}}$ is either $(\epsilon,{\rm EC}1)$-extremal or $(\epsilon,{\rm EC}2)$-extremal, and $D_{c_{2}}$ is $(\epsilon,{\rm EC}3)$-extremal. Hence $x\in A_{c_{1}}\cup B_{c_{1}}$ and $y\in W_{c_{2}}^{1}\cup W_{c_{2}}^{3}$. Choose $z\in N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)$. Note that $D_{c_{1}}$ and $D_{c_{2}}$ are $\delta$-crossing. Then $|X_{c_{1}}\cap Y_{c_{2}}|\geq\frac{\delta n}{4}$ whenever $X\in\{A,B\}$ and $Y\in\{W^{1},W^{3}\}$. By Lemma 3.2, there are $Z\in\{A,B\}$ and $W\in\{W^{2},W^{4}\}$ such that $d_{{c_{1}}}^{+}(x,Z_{c_{1}})\geq|Z_{c_{1}}|-\epsilon n$ and $d_{{c_{2}}}^{-}(y,W_{c_{2}})\geq|W_{c_{2}}|-2\epsilon n$. If $|Z_{c_{1}}\cap W_{c_{2}}|\geq\frac{\delta n}{4}$, then $|N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)|\geq|Z_{c_{1}}\cap W_{c_{2}}|-3\epsilon n\geq\frac{\delta n}{5}$, as desired.
Now, we consider $|Z_{c_{1}}\cap W_{c_{2}}|<\frac{\delta n}{4}$. Without loss of generality, assume that $Z=A$ and $W=W^{2}$. It follows that $|A_{c_{1}}\cap W_{c_{2}}^{4}|\geq|W_{c_{2}}^{4}|-(\frac{\delta}{4}+2\epsilon)n$. This implies $|A_{c_{1}}\cap W_{c_{2}}^{3}|\geq|C_{c_{2}}^{4}|-(\frac{\delta}{4}+2\epsilon)n\geq\frac{\delta n}{2}$ since $|C^{2}_{4}|\geq\delta n$. If $|W_{c_{2}}^{1}\cap N_{{c_{4}}}^{-}(v_{2}^{1})|\geq\frac{\delta n}{9}$, then choose $y\in W_{c_{2}}^{1}\cap N_{{c_{4}}}^{-}(v_{2}^{1})$. Hence $|N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)|\geq|A_{c_{1}}\cap W_{c_{2}}^{4}|-3\epsilon n\geq\frac{\delta n}{5}$. If $|W_{c_{2}}^{1}\cap N_{{c_{4}}}^{-}(v_{2}^{1})|<\frac{\delta n}{9}$, then $|W_{c_{2}}^{3}\cap N_{{c_{4}}}^{-}(v_{2}^{1})|\geq|W_{c_{2}}^{3}|-(\frac{\delta}{9}+2\epsilon)n$. Thus, $|N_{{c_{1}}}^{+}(x)\cap N_{{c_{4}}}^{-}(v_{2}^{1})|\geq|A_{c_{1}}\cap W_{c_{2}}^{3}|-(3\epsilon+\frac{\delta}{9})n\geq\frac{\delta n}{9}$. Now, we connect $P^{1}$ and $P^{2}$ by using a rainbow directed path $P^{1}v_{4}^{1}xyv_{1}^{2}P^{2}$ with colors ${\rm col}(P^{1}\cup P^{2})\cup\{c_{1},c_{3},c_{4}\}$, as desired.
@@
In view of Claim 3.11, there are at least $\frac{\delta n}{10}$ choices for each of $c_{1},c_{2},c_{3},c_{4},x,y,(z)$ given any previous choices. Thus, we obtain the desired rainbow directed path $P^{1}v_{4}^{1}xzyv_{1}^{2}P^{2}$ or $P^{1}v_{4}^{1}xyv_{1}^{2}P^{2}$. Applying the above process for each pair $(P^{j},P^{j+1})$ with $j\in[s]$ where $P^{s+1}:=P^{1}$. Every time we use at most four unused colors in $\overline{\mathcal{U}}$ and at most three unused vertices. Therefore, for each choice, at most $4s\leq 4t\leq 8\kappa n\ll\frac{\delta n}{10}$ colors are forbidden and $3s\ll\frac{\delta n}{10}$ vertices are forbidden. Thus we obtain a rainbow directed cycle $C$ of length at most $7s$ which contains every $P^{j}$ with $j\in[s]$ as a segment.